# AI-Native 6G Physical Layer with Cross-Module Optimization and Cooperative Control Agents

### v1

@@ I Introduction
The generational evolution of wireless communication systems is inherently intertwined with the deep integration of transformative technologies, especially a series of innovations in the physical layer [1, 2, 3]. Specifically, due to the successful application of artificial intelligence (AI) in the field of computer vision and natural language processing, the enhancement of wireless communication using AI has attracted great attention in recent years [4], where the data-driven nature and nonlinear processing capabilities of AI solutions bring more adaptability and performance gains compared to other traditional solutions. A series of creative academic works have been done one after another, such as channel state information (CSI) feedback [5, 6, 7, 8], channel estimation [9, 10], beamforming [11], modulation [12, 13] and so on. The robustness of channel estimation under imperfect CSI conditions has been further extensively studied in the fifth generation (5G) systems [14], while emerging The sixth generation (6G) technologies are pushing the boundaries of ultra-reliable low-latency communications through novel physical-layer innovations [15]. For the standardization process of 5G-Advanced, the 3rd Generation Partnership Project (3GPP) has also initiated studies on AI/machine learning (ML) for new radio (NR) air interface enhancements [16], focusing on three pivotal use cases: CSI feedback enhancement, beam management, and high-precision positioning. These studies encompass evaluation methodologies, potential specification impacts, and other relevant aspects. Specifically, regarding CSI feedback enhancement, 3GPP Release 18 and Release 19 standardization efforts have centered on two technical directions: AI/ML-based CSI compression and AI/ML-based user equipment (UE)-side CSI prediction. These efforts aim to surpass the performance limitations of conventional codebook-based CSI feedback mechanisms [17, 18, 19, 20] while simultaneously reducing overhead and latency. For AI/ML-based beam management, the standardization focuses on spatial and temporal domain beam prediction to achieve reduced reference signal overhead and real-time beam information acquisition in high frequency bands. As for high-precision positioning, 3GPP aims to leverage AI/ML to significantly improve localization accuracy in non-line-of-sight (NLoS) dominated environments.
@@
However, the above works in academia and industry still adopt a modular architecture, which decouples the physical layer into separate functional blocks and optimizing them independently. Although this compartmentalized design philosophy is beneficial for finding the optimal solution of independent modules and achieve engineering tractability, its inherent limitation, i.e., the globally suboptimal performance caused by module isolation, has become increasingly evident.
@@
Fortunately, with the introduction of AI/ML solutions, holistic design beyond existing modular architectures has become possible. The 6G presents an opportunity to revisit the fundamentals of radio system design and explore native integration with AI/ML, targeting transformative leaps in capability, efficiency and simplicity, which requires a holistic and optimized design across all layers of the radio interface and architecture [21, 22, 23, 24, 25, 26, 27, 28]. Related works demonstrate a paradigm shift in research focus from module-level AI/ML features to systematic joint design. At the transmitter, joint design of the modulation and precoding [29], joint source-channel coding (JSCC) [30, 31, 32] and modulation [33, 34] are studied. At the receiver, [35] considers the joint design of channel estimation and data detection, realizing an integrated receiver by end-to-end training procedure. As for the downlink transmission with uplink feedback procedure, existing works propose joint CSI feedback and precoding [36, 37, 38] with pilot design [39, 40] or channel estimation [41, 42].
@@
However, although the existing works achieve joint design of partial modules of the link, they still fail to consider more thorough liberation of the global freedom degree for design of physical layer with AI/ML features, i.e., i) simultaneous optimization across global modules of uplink AI/ML-based joint channel-source coding with modulation, and downlink AI/ML-based modulation with precoding and corresponding data detection, and ii) joint exploitation of the features across transmitting layers to expand the solution space for finding the optimal constellation and precoder. Moreover, existing works largely rely on idealized system assumptions, e.g., additive white Gaussian noise (AWGN) downlink channel or ideal uplink feedback, may encountering intractable challenges of adaptability when scaling to practical scenarios involving more practical channel model and imperfect uplink channel for CSI feedback.
@@
  • Considering current state-of-the-art study, the model lifecycle management (LCM) [43] in 3GPP studies alleviates the above difficulties through performance monitoring, model switching, and model updating procedure. However, it still poses a series of challenges. First, existing LCM solutions employ post-hoc adaptation with lower real-time performance, as switching is triggered only after performance degradation is detected during model monitoring. Besides, since more and more features with two-sided models are anticipated to be integrated into 6G systems, the existing LCM solutions will inevitably incur substantial signaling overhead and impose heavy standardization burdens, where each new two-sided feature requires additional control signaling over the air interface to ensure model alignment.
@@
  • An AI-native framework is proposed, which involves i) AI/ML-based control strategy capable of dynamically selecting the optimal transmission scheme based on real-time wireless channel conditions, while establishing two-sided model alignment between UE and NW, and ii) simultaneous optimization across global modules of uplink AI/ML-based joint channel-source coding with modulation, and downlink AI/ML-based modulation with precoding and corresponding data detection, breaking down traditional inter-module information silos to enable end-to-end optimization under global objectives.
@@
Throughout this paper, upper-case and lower-case letters denote scalars. Boldface upper-case and boldface lower-case letters denote matrices and vectors, respectively. Calligraphic upper-case letters denote sets. Specifically, the set of real and complex numbers are denoted by $\mathbb{R}$ and $\mathbb{C}$, respectively. To further enhance the readability and accessibility of the manuscript, a table summarizing the key mathematical notations with their definitions can be find in Table I.
@@ II-A MIMO-OFDM Systems
A typical single-user MIMO-OFDM system operating in frequency division duplexing (FDD) mode with $N_{\mathrm{tx}}$ transmit antennas at base station (BS) and $N_{\mathrm{rx}}$ receive antennas at UE is assumed. Note that while time division duplexing (TDD) systems with reciprocity-based beamforming offer advantages in massive MIMO scenarios by eliminating downlink training overhead, the FDD framework is adopted in this work due to its compatibility with dominant 5G NR deployments. The proposed AI-native enhancements to CSI feedback remain relevant even in hybrid TDD/FDD systems where feedback is still required. The study considers signal transmission based on OFDM frame structure with $N_{\mathrm{sc}}$ subcarriers and $N_{\mathrm{t}}$ OFDM symbols, where the subcarrier index and OFDM symbol index are represent by $f\in\{1,\ldots,N_{\mathrm{sc}}\}$ and $t\in\{1,\ldots,N_{\mathrm{t}}\}$, respectively. Given a sufficiently long cyclic prefix to mitigate inter-symbol interference, the baseband-equivalent MIMO transmission process for each temporal-frequency resource element (RE) in the OFDM grid can be expressed as,
@@
where $\mathbf{y}_{f,t}\in\mathbb{C}^{N_{\mathrm{rx}}\times 1}$ is the received signal, $\mathbf{H}_{f,t}\in\mathbb{C}^{N_{\mathrm{rx}}\times N_{\mathrm{tx}}}$ is the downlink channel matrix, $\mathbf{x}_{f,t}\in\mathbb{C}^{N_{\mathrm{tx}}\times 1}$ is the precoded transmit signal and $\mathbf{n}_{f,t}\in\mathbb{C}^{N_{\mathrm{rx}}\times 1}$ is the additive white Gaussian noise with variance $\sigma_{n}^{2}$. Correspondingly, $\mathbf{y}\in\mathbb{C}^{N_{\mathrm{rx}}\times N_{\mathrm{sc}}\times N_{\mathrm{t}}}$, $\mathbf{H}\in\mathbb{C}^{N_{\mathrm{rx}}\times N_{\mathrm{tx}}\times N_{\mathrm{sc}}\times N_{\mathrm{t}}}$, $\mathbf{x}\in\mathbb{C}^{N_{\mathrm{tx}}\times N_{\mathrm{sc}}\times N_{\mathrm{t}}}$ and $\mathbf{n}\in\mathbb{C}^{N_{\mathrm{rx}}\times N_{\mathrm{sc}}\times N_{\mathrm{t}}}$ represents the received signal, downlink channel, precoded transmit signal and noise on all subcarriers and OFDM symbols, respectively. Based on the above system, data transmission and CSI feedback via downlink and uplink channel respectively can be achieved.
@@ II-B Data Transmission via Downlink Channel
At the BS side, the downlink processing starts with the encoding of information bit stream $\mathbf{b}\in\{0,1\}^{N_{\mathrm{b}}\times 1}$ using a 5G-compliant low-density parity-check (LDPC) channel encoder, resulting in encoded bit stream $\mathbf{c}\in\{0,1\}^{N_{\mathrm{c}}\times 1}$. The coded bit stream undergoes modulation to form complex symbols $\mathbf{s}\in\widetilde{\mathbb{C}}^{N_{\mathrm{s}}\times 1}$, where $\widetilde{\mathbb{C}}$ denotes the constellation set and the symbol sequence length satisfies $N_{\mathrm{s}}=N_{\mathrm{c}}/m$ with $m$ denoting the modulation order. The modulated symbols with inserted demodulation reference signals (DMRS), undergo precoding through multiplication with a precoding matrix $\mathbf{P}\in\mathbb{C}^{N_{\mathrm{tx}}\times N_{\mathrm{layer}}}$ to form the transmit signal $\mathbf{x}$. The selection of $N_{\mathrm{layer}}$ and corresponding $\mathbf{P}$ can be dynamically adapted to channel conditions. To mitigate CSI feedback overhead, the whole band of $N_{\mathrm{sc}}$ subcarriers are uniformly divided into $N_{\mathrm{sb}}$ subbands, with all symbols corresponding to the same subband sharing a common precoding matrix. The subband index and layer index are represented by $k\in\{1,\ldots,N_{\mathrm{sb}}\}$ and $n_{\mathrm{layer}}\in\{1,\ldots,N_{\mathrm{layer}}\}$, respectively. Subsequently, the precoded symbols are transformed into OFDM waveform through an inverse fast Fourier transform (IFFT) operation followed by cyclic prefix (CP) insertion.
@@ II-C CSI Feedback via Uplink Channel
where $\mathbf{\Lambda}_{k}=\mathrm{diag}\{\lambda_{k,1},\lambda_{k,2},\ldots,\lambda_{k,N_{\mathrm{tx}}}\}$ is a diagonal matrix of eigenvalues sorted in descending order. The top $N_{\mathrm{layer}}$ eigenvectors, concatenated and denoted as $\mathbf{w}_{k}\in\mathbb{C}^{N_{\mathrm{tx}}N_{\mathrm{layer}}\times 1}$, will be extracted and therefore the CSI matrix to be feedback can be written as,
$\mathbf{W}=[\mathbf{w}_{1},\mathbf{w}_{2},\cdots,\mathbf{w}_{N{\mathrm{sb}}}]^{\mathrm{T}}\in\mathbb{C}^{N_{\mathrm{sb}}\times N_{\mathrm{tx}}N_{\mathrm{layer}}}$ (5)
With the preprocessed CSI available, the UE compresses it and feeds it back to the BS via the noisy uplink channel, enabling the BS to reconstruct the CSI for downlink precoding. In 5G systems, the standardized Type I and enhanced Type II (eType II) codebooks are widely used for CSI feedback [17, 18, 19, 20], where the UE selects precoding matrix from predefined codebooks and report the precoding matrix index (PMI), allowing the BS to reconstruct the CSI accordingly. Additionally, AI/ML-based CSI feedback has been discussed in 3GPP, employing paired encoder-decoder NNs $f_{\rm e}(\cdot;\Theta_{\rm{E}})$ and $f_{\rm d}(\cdot;\Theta_{\rm{D}})$ with trainable parameters $\Theta_{\rm{E}}$ and $\Theta_{\rm{D}}$ at the UE and BS, respectively. Thus the AI/ML-based autoencoder $f_{\rm a}(\cdot;\Theta_{\rm{A}})$ with trainable parameters $\Theta=\{\Theta_{\rm{E}},\Theta_{\rm{D}}\}$ for CSI feedback can be represented as
@@
where $f_{\rm e}(\cdot;\Theta_{\rm{E}})$ and $f_{\rm d}(\cdot;\Theta_{\rm{D}})$ implement the CSI compression and reconstruction processes respectively. Note that $n_{\rm{channel}}(\cdot)$ denotes the practical noisy uplink channel, where the decoder at BS tolerates the lossy input information and introduces disturbances to the decoder output, which can be comparable to and expended to the scenarios with disturbances caused by CSI errors introduced at the UE side.
@@
III AI-Native Cross-Module Optimized Physical Layer with Cooperated Control Agent

### v2

@@ I Introduction
The generational evolution of wireless communication systems is inherently intertwined with the deep integration of transformative technologies, especially a series of innovations in the physical layer series2015imt ; 5Ghai ; TS38913 . Specifically, due to the successful application of artificial intelligence (AI) in the field of computer vision and natural language processing, the enhancement of wireless communication using AI has attracted great attention in recent years huang2019deep , where the data-driven nature and nonlinear processing capabilities of AI solutions bring more adaptability and performance gains compared to other traditional solutions. A series of creative academic works have been done one after another, such as channel state information (CSI) feedback han2024ai ; xiao2021ai ; liu2021evcsinet ; bi2022novel , channel estimation le2021deep ; gao2023deep , beamforming al2022review , modulation alberge2018deep ; jiang2019deep and so on. The robustness of channel estimation under imperfect CSI conditions has been further extensively studied in the fifth generation (5G) systems pourkabirian2021robust , while emerging sixth generation (6G) technologies are pushing the boundaries of ultra-reliable low-latency communications through novel physical-layer innovations pourkabirian2024vision . For the standardization process of 5G-Advanced, the 3rd Generation Partnership Project (3GPP) has also initiated studies on AI/machine learning (ML) for new radio (NR) air interface enhancements 213599 , focusing on three pivotal use cases: CSI feedback enhancement, beam management, and high-precision positioning. These studies encompass evaluation methodologies, potential specification impacts, and other relevant aspects. Specifically, regarding CSI feedback enhancement, 3GPP Release 18 and Release 19 standardization efforts have centered on two technical directions: AI/ML-based CSI compression and AI/ML-based user equipment (UE)-side CSI prediction. These efforts aim to surpass the performance limitations of conventional codebook-based CSI feedback mechanisms TS0000 ; TS0001 ; TS0002 ; TS0003 while simultaneously reducing overhead and latency. For AI/ML-based beam management, the standardization focuses on spatial and temporal domain beam prediction to achieve reduced reference signal overhead and real-time beam information acquisition in high frequency bands. As for high-precision positioning, 3GPP aims to leverage AI/ML to significantly improve localization accuracy in non-line-of-sight (NLoS)-dominated environments.
@@
However, the above works in academia and industry still adopt a modular architecture, which decouples the physical layer into separate functional blocks and optimizes them independently. Although this compartmentalized design philosophy is beneficial for finding the optimal solution of independent modules and achieving engineering tractability, its inherent limitation, i.e., the globally suboptimal performance caused by module isolation, has become increasingly evident.
@@
Fortunately, with the introduction of AI/ML solutions, holistic design beyond existing modular architectures has become possible. The 6G presents an opportunity to revisit the fundamentals of radio system design and explore native integration with AI/ML, targeting transformative leaps in capability, efficiency and simplicity, which requires a holistic and optimized design across all layers of the radio interface and architecture 251881 ; jiang2021road ; yang20226g ; hoydis2021toward ; jiao2024advanced ; ozpoyraz2022deep ; abd2024evolution ; salh2021survey . Related works demonstrate a paradigm shift in research focus from module-level AI/ML features to systematic joint design. At the transmitter, joint design of the modulation and precoding zhang2023multi , joint source-channel coding (JSCC) xu2022deep ; chu2023deep ; jiang2024deep and modulation matsumoto2023impact ; inoue2021deep are studied. At the receiver, ait2021end considers the joint design of channel estimation and data detection, realizing an integrated receiver by end-to-end training procedure. As for the downlink transmission with uplink feedback procedure, existing works propose joint CSI feedback and precoding sun2022deep ; sohrabi2021deep ; guo2024deep with pilot design jang2022deep ; jang2019deep or channel estimation guo2025deep ; zhao2022joint .
@@
However, although the existing works achieve joint design of partial modules of the link, they still fail to consider more thorough exploration of the global degrees of freedom for design of physical layer with AI/ML features, i.e., i) simultaneous optimization across global modules of uplink AI/ML-based joint source-channel coding with modulation, and downlink AI/ML-based modulation with precoding and corresponding data detection, and ii) joint exploitation of the features across transmitting layers to expand the solution space for finding the optimal constellation and precoder. Moreover, existing works largely rely on idealized system assumptions, e.g., additive white Gaussian noise (AWGN) downlink channel or ideal uplink feedback, may encounter intractable challenges of adaptability when scaling to practical scenarios involving more practical channel model and imperfect uplink channel for CSI feedback.
@@
  • Considering current state-of-the-art study, the model lifecycle management (LCM) chen20235g in 3GPP studies alleviates the above difficulties through performance monitoring, model switching, and model updating procedure. However, it still poses a series of challenges. First, existing LCM solutions employ post-hoc adaptation with lower real-time performance, as switching is triggered only after performance degradation is detected during model monitoring. Besides, since more and more features with two-sided models are anticipated to be integrated into 6G systems, the existing LCM solutions will inevitably incur substantial signaling overhead and impose heavy standardization burdens, where each new two-sided feature requires additional control signaling over the air interface to ensure model alignment.
@@
  • An AI-native framework is proposed, which involves i) AI/ML-based control strategy capable of dynamically selecting the optimal transmission scheme based on real-time wireless channel conditions, while establishing two-sided model alignment between UE and NW, and ii) simultaneous optimization across global modules of uplink AI/ML-based joint source-channel coding with modulation, and downlink AI/ML-based modulation with precoding and corresponding data detection, breaking down traditional inter-module information silos to enable end-to-end optimization under global objectives.
@@
Throughout this paper, upper-case and lower-case letters denote scalars. Boldface upper-case and boldface lower-case letters denote matrices and vectors, respectively. Calligraphic upper-case letters denote sets. Specifically, the sets of real and complex numbers are denoted by $\mathbb{R}$ and $\mathbb{C}$, respectively. To further enhance the readability and accessibility of the manuscript, a table summarizing the key mathematical notations with their definitions can be found in Table I.
@@ II-A MIMO-OFDM Systems
A typical single-user MIMO-OFDM system operating in frequency division duplexing (FDD) mode with $N_{\mathrm{tx}}$ transmit antennas at base station (BS) and $N_{\mathrm{rx}}$ receive antennas at UE is assumed. Note that while time division duplexing (TDD) systems with reciprocity-based beamforming offer advantages in massive MIMO scenarios by eliminating downlink training overhead, the FDD framework is adopted in this work due to its compatibility with dominant 5G NR deployments. The proposed AI-native enhancements to CSI feedback remain relevant even in hybrid TDD/FDD systems where feedback is still required. The study considers signal transmission based on OFDM frame structure with $N_{\mathrm{sc}}$ subcarriers and $N_{\mathrm{t}}$ OFDM symbols, where the subcarrier index and OFDM symbol index are represented by $f\in\{1,\ldots,N_{\mathrm{sc}}\}$ and $t\in\{1,\ldots,N_{\mathrm{t}}\}$, respectively. Given a sufficiently long cyclic prefix to mitigate inter-symbol interference, the baseband-equivalent MIMO transmission process for each temporal-frequency resource element (RE) in the OFDM grid can be expressed as,
@@
where $\mathbf{y}_{f,t}\in\mathbb{C}^{N_{\mathrm{rx}}\times 1}$ is the received signal, $\mathbf{H}_{f,t}\in\mathbb{C}^{N_{\mathrm{rx}}\times N_{\mathrm{tx}}}$ is the downlink channel matrix, $\mathbf{x}_{f,t}\in\mathbb{C}^{N_{\mathrm{tx}}\times 1}$ is the precoded transmit signal and $\mathbf{n}_{f,t}\in\mathbb{C}^{N_{\mathrm{rx}}\times 1}$ is the additive white Gaussian noise with variance $\sigma_{n}^{2}$. Correspondingly, $\mathbf{y}\in\mathbb{C}^{N_{\mathrm{rx}}\times N_{\mathrm{sc}}\times N_{\mathrm{t}}}$, $\mathbf{H}\in\mathbb{C}^{N_{\mathrm{rx}}\times N_{\mathrm{tx}}\times N_{\mathrm{sc}}\times N_{\mathrm{t}}}$, $\mathbf{x}\in\mathbb{C}^{N_{\mathrm{tx}}\times N_{\mathrm{sc}}\times N_{\mathrm{t}}}$ and $\mathbf{n}\in\mathbb{C}^{N_{\mathrm{rx}}\times N_{\mathrm{sc}}\times N_{\mathrm{t}}}$ represent the received signal, downlink channel, precoded transmit signal and noise on all subcarriers and OFDM symbols, respectively. Based on the above system, data transmission and CSI feedback via downlink and uplink channel respectively can be achieved.
@@ II-B Data Transmission via Downlink Channel
At the BS side, the downlink processing starts with the encoding of information bit stream $\mathbf{b}\in\{0,1\}^{N_{\mathrm{b}}\times 1}$ using a 5G-compliant low-density parity-check (LDPC) channel encoder, resulting in encoded bit stream $\mathbf{c}\in\{0,1\}^{N_{\mathrm{c}}\times 1}$. The coded bit stream undergoes modulation to form complex symbols $\mathbf{s}\in\widetilde{\mathbb{C}}^{N_{\mathrm{s}}\times 1}$, where $\widetilde{\mathbb{C}}$ denotes the constellation set and the symbol sequence length satisfies $N_{\mathrm{s}}=N_{\mathrm{c}}/m$ with $m$ denoting the modulation order. The modulated symbols with inserted demodulation reference signals (DMRS), undergo precoding through multiplication with a precoding matrix $\mathbf{P}\in\mathbb{C}^{N_{\mathrm{tx}}\times N_{\mathrm{layer}}}$ to form the transmit signal $\mathbf{x}$. The selection of $N_{\mathrm{layer}}$ and corresponding $\mathbf{P}$ can be dynamically adapted to channel conditions. To mitigate CSI feedback overhead, the whole band of $N_{\mathrm{sc}}$ subcarriers is uniformly divided into $N_{\mathrm{sb}}$ subbands, with all symbols corresponding to the same subband sharing a common precoding matrix. The subband index and layer index are represented by $k\in\{1,\ldots,N_{\mathrm{sb}}\}$ and $n_{\mathrm{layer}}\in\{1,\ldots,N_{\mathrm{layer}}\}$, respectively. Subsequently, the precoded symbols are transformed into OFDM waveform through an inverse fast Fourier transform (IFFT) operation followed by cyclic prefix (CP) insertion.
@@ II-C CSI Feedback via Uplink Channel
where $\mathbf{\Lambda}_{k}=\mathrm{diag}\{\lambda_{k,1},\lambda_{k,2},\ldots,\lambda_{k,N_{\mathrm{tx}}}\}$ is a diagonal matrix of eigenvalues sorted in descending order. The top $N_{\mathrm{layer}}$ eigenvectors, concatenated and denoted as $\mathbf{w}_{k}\in\mathbb{C}^{N_{\mathrm{tx}}N_{\mathrm{layer}}\times 1}$, will be extracted, and therefore the CSI matrix to be fed back can be written as,
$\mathbf{W}=[\mathbf{w}_{1},\mathbf{w}_{2},\cdots,\mathbf{w}_{N_{\mathrm{sb}}}]^{\mathrm{T}}\in\mathbb{C}^{N_{\mathrm{sb}}\times N_{\mathrm{tx}}N_{\mathrm{layer}}}$ (5)
With the preprocessed CSI available, the UE compresses it and feeds it back to the BS via the noisy uplink channel, enabling the BS to reconstruct the CSI for downlink precoding. In 5G systems, the standardized Type I and enhanced Type II (eType II) codebooks are widely used for CSI feedback TS0000 ; TS0001 ; TS0002 ; TS0003 , where the UE selects a precoding matrix from predefined codebooks and reports the precoding matrix index (PMI), allowing the BS to reconstruct the CSI accordingly. Additionally, AI/ML-based CSI feedback has been discussed in 3GPP, employing paired encoder-decoder NNs $f_{\rm e}(\cdot;\Theta_{\rm{E}})$ and $f_{\rm d}(\cdot;\Theta_{\rm{D}})$ with trainable parameters $\Theta_{\rm{E}}$ and $\Theta_{\rm{D}}$ at the UE and BS, respectively. Thus the AI/ML-based autoencoder $f_{\rm a}(\cdot;\Theta_{\rm{A}})$ with trainable parameters $\Theta=\{\Theta_{\rm{E}},\Theta_{\rm{D}}\}$ for CSI feedback can be represented as
@@
where $f_{\rm e}(\cdot;\Theta_{\rm{E}})$ and $f_{\rm d}(\cdot;\Theta_{\rm{D}})$ implement the CSI compression and reconstruction processes respectively. Note that $n_{\rm{channel}}(\cdot)$ denotes the practical noisy uplink channel, where the decoder at BS tolerates the lossy input information and introduces disturbances to the decoder output, which can be comparable to and extended to the scenarios with disturbances caused by CSI errors introduced at the UE side.
@@
III AI-Native Cross-Module Optimized Physical Layer with Cooperative Control Agents